OpenStudy (anonymous):

For the given statement Pn, write the statements P1, Pk, and Pk+1. 2 + 4 + 6 + . . . + 2n = n(n+1)

OpenStudy (anonymous):

pk+1 is included in the statement.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

3. P1: 2=1(1+1) Pk: 2 + 4 + 6 + . . . + 2k = k(k+1) Pk+1: 2 + 4 + 6 + . . . + 2(k+1) = (k+1)[(k+1)+1)]